OpenStudy (anonymous):

What is 5.6 × 10−3 yd/year in feet per day? A. 6.813 × 10−7 ft/day B. 6.132 × 10−6 ft/day C. 5 × 10−6 ft/day D. 4.6 × 10−5 ft/day

ganeshie8 (ganeshie8):

3 feet = 1 yard 365 days = 1 year

ganeshie8 (ganeshie8):

\[\large 5.6 × 10^{−3} \dfrac{\text{yard}}{\text{year}} \times \dfrac{\text{3 feet}}{\text{1 yard}} \times \dfrac{\text{1 year}}{\text{365 days}} \]

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

\[5.6\times10^{-3}~yd/yr=\frac{ 5.6\times10^{-3}\times3 }{ 365 }ft~/day\]

OpenStudy (anonymous):

what is the answer though ? D? @ganeshie8

OpenStudy (trinibutera):

@Wonderland_11 mhm